# A Liquid-Core Fiber Platform for Classical and Entangled Two-Photon Absorption Measurements

### v1

@@ Abstract
We introduce a toluene-filled fiber platform for two-photon absorption measurements. By confining both the light and molecular sample inside the 5 $\mu$m hollow core of the fiber, we increase the distance over which the nonlinear light-matter interaction occurs. With only a 7.3 nL excitation volume, we measure classical two-photon absorption (C2PA) at an average laser power as low as 1.75 nW, which is a 45-fold improvement over a conventional free-space technique. We use this platform to attempt to measure entangled two-photon absorption (E2PA), a process with a limited operating regime due to a crossover in dominating processes from E2PA to C2PA as photon flux is increased. Recently, several teams of researchers have reported that E2PA cross sections are much smaller than previously claimed. As a result, the process dominates at photon fluxes so low that it is extremely difficult or impossible to measure using conventional free-space techniques. In this report, we implement the first E2PA measurement using a waveguide. We see no evidence of E2PA, and we set an upper bound on the cross section consistent with these recent reports.
@@
1 Introduction
@@
Here, we show that with a toluene-filled fiber, 2PA cross sections, which are the quantities that describe the strength of the 2PA process, can be measured at extremely low powers. To the best of our knowledge, this work is the first to show a 2PA measurement at an average power as low as 1.75 nW. This is a 45-fold lower power than that achieved in our previous results16 for low-power 2PA in free space for the same sample. These fiber and free-space measurements serve as a one-to-one comparison since both were designed to test the sensitivity limits of the platforms. We believe this is the first time that the advantage of the fiber platform for 2PA has been quantified.
@@
This sensitivity, paired with the ability to increase our laser power by up to nine orders of magnitude, demonstrates that the platform is well equipped to boost previously undetectable signals. This is necessary, for example, for measurements of C2PA cross sections at much less efficient wavelengths or for much less efficient absorbers. Furthermore, rather than being limited to a high peak power laser17, unconventional light sources may produce sizeable 2PA signals.
@@
One such light source is an entangled photon pair source based on spontaneous parametric downconversion (SPDC). We employ an SPDC source in an attempt to measure entangled two-photon absorption (E2PA). To differentiate between E2PA and 2PA excited by a laser source, we refer to the latter process as classical two-photon absorption (C2PA). The E2PA process has been a subject of theoretical study since the 1980s (Refs. 18, 19, 20), driven by the idea that the strong temporal and spatial correlations of SPDC photons are ideal for 2PA. These temporal and spatial correlations are characterized by the entanglement time and area, respectively20. This absorption process is known to scale linearly with the excitation photon flux, rather than the quadratic scaling of C2PA, at low photon flux where it dominates over C2PA and thus demonstrates a quantum advantage.
@@
Recently, a wide range of evidence from a variety of experimental research studies has shown that molecular E2PA cross sections are much smaller than previously reported. The majority of studies have reported no E2PA signal above the noise floor (Refs. 21, 16, 22, 23, 24, 25, 26, 27, 28, 29, 30). One study has reported an extremely weak signal (Ref. 31). These experimental reports have been reinforced by recent theoretical works that show that the linear scaling will dominate at photon fluxes so low that the resulting minuscule signal will be difficult or impossible to measure with current technology (Refs. 32, 33, 34, 35). Furthermore, the conditions under which the minuscule signal is maximized require careful selection of the SPDC source and two-photon absorber. These experimental and theoretical studies all consider conventional free-space excitation conditions. Here we present, to the best of our knowledge, the first E2PA measurement implemented in a waveguide. We do not observe E2PA and set an upper bound on the E2PA cross section of the molecule AF455.
@@
In this paper, we discuss the principles of our experimental scheme and the design considerations along with our experimental implementation and its characterization. Next, we present our results for excitation of the molecule AF455 with both laser and SPDC photons and derive the C2PA and E2PA cross sections, respectively. Finally we compare our results with those from previous free-space measurements16. More details on experimental characterization, data acquisition and data analysis are provided in the supplemental information.
2 Operating Principles
@@
The benefit of this scheme is illustrated in Fig. 1(a-b). In the free-space approach (Fig. 1(a)), light is focused into the cuvette to a waist $w_{0}$ and diverges on a length scale characterized by the Rayleigh range $z_{R}$ which scales quadratically with $w_{0}$. Thus, increasing beam intensity occurs at the cost of maintaining the intensity over a shorter distance $z_{R}$. For typical focusing conditions of $w_{0}\approx 30\,\mu$m for an 810 nm beam into toluene, $z_{R}\approx 5.2$ mm, which is about half of the cuvette’s length. In contrast, with the fiber-based approach (Fig. 1(b)), the beam waist $w_{0}$ is maintained with low loss over the entire propagation length of the fiber because of optical confinement. Typically, $w_{0}$ is much smaller in fiber than in free space. This high light intensity is advantageous for 2PA because two photons must be spatially and temporally overlapped at a molecule for absorption to occur. Maintaining the peak intensity of the focused beam over centimeter length-scales increases the excitation probability because of the interaction with a large number of molecules.
@@
The refractive indices as a function of wavelength for the toluene core and silica cladding are plotted in Fig. 1(c) using the known dispersion equations for the materials36, 37. The effective index of refraction of the fundamental and second-order modes are shown and are calculated using Ref. 38. We can estimate the number of modes that can propagate along the fiber, which has a core diameter $d$, using the V-number,
@@
In addition to toluene allowing broadband guidance of light in a silica fiber, this solvent also has a low absorption coefficient39 at the excitation and fluorescence wavelengths (0.0030 cm-1 and 0.0039 cm-1, respectively). This is a necessary condition otherwise the long length of the fiber will add little to no benefit since all the light would be absorbed after a short distance. For comparison, water has an absorption coefficient that is about one order of magnitude larger at 810 nm (0.0209 cm-139).
3 Materials and Methods
To prepare the fiber, we first cut clean facets on both ends to minimize light-coupling losses. We use a custom-built coil heater to remove about 20 mm of the polyamide coating from each end, then we cleave the fiber ends in the regions where the coatings were removed. We inspect the ends under a digital microscope (Keyence VHX 7000) to ensure the cuts are smooth and to check for particle contamination. Typical images are shown in Fig. 2(a)-(b). The fiber ends are placed in a tubing sleeve to prevent breakage and to allow for simple attachment to the custom-built fiber adapters (similar to those used in Refs. 40, 41, technical drawings provided in Ref. 42).
Figure 2: Digital microscope images of one end of (a) the fiber and (b) the fiber inside of the tubing sleeve. (c) sCMOS image of the 810 nm excitation laser guided through the fiber. (d) EMCCD image of the fluorescence from AF455 guided out of the fiber. The last two images are taken in the image plane of the fiber face. Some features of the tube and tubing sleeve are visible in both images as described in Supporting Section 6.
@@ 3 Materials and Methods
The optical and microfluidic setup is illustrated in Fig. 3. Here we give a brief overview of the setup. A complete list of the components is given in Ref. 42. A tunable femtosecond pulsed laser operated at 810 nm (110 fs pulses) is used to generate SPDC photons for E2PA measurements and its direct output is used for C2PA measurements. For photon pair generation, the output is first directed through a second harmonic generation unit to produce 405 nm light. The average power is controlled using a half-wave plate and a polarizer. Spectral filters are used to remove any remaining 810 nm light or unwanted harmonics. A small portion of this beam is picked off by a beam sampler to monitor the power and beam pointing stability of the laser on a quadrant (quad) detector. The blue light is focused ($f=300$ mm lens) into a 10 mm-long type-0 periodically-poled potassium titanyl phosphate (ppKTP) crystal to generate photon pairs centered at 810 nm. The crystal is temperature controlled at $40.00\pm 0.01\,^{\circ}$. The photon pairs are collimated with a 50 mm lens and any remaining pump light is filtered out.
@@
We measure that this source produces $8.25\times 10^{9}$ photons s-1 mW-1 using a scientific Complementary Metal–Oxide–Semiconductor (sCMOS) camera placed directly after the crystal and collimating lens. In our previous report16 we found that this source has a spectral bandwith of approximately 75 nm full width at half maximum (FWHM) under these operating conditions. The entanglement time of the SPDC is estimated using a discrete Fourier transform of the joint spectral amplitude, which is estimated using the measured joint spectral intensity (shown in Ref. 16) and a phase estimated using the calculated group delay dispersion accumulated by the SPDC between the center of the crystal and its position along the fiber length. The broadening of the entanglement time across the length of the fiber is taken into account in our calculations (Supporting Section 10). At the entrance of the fiber the entanglement time is 1070 fs. The entanglement area was not measured, but we estimate it to be in the range of 2.1 $\mu$m2 to 18 $\mu$m2. The lower bound is set by a diffraction limited spot size (see Ref. 16). The upper bound is set by the spot size of the fundamental mode of the fiber.
@@
For the C2PA measurements, the pulsed laser is directed around the nonlinear crystals and through a half-wave plate, polarizing beamsplitter and neutral density filters for control over the power. The light from both sources is recombined at a back-polished mirror. Either source is directed into the fiber-based two-photon excited fluorescence (2PEF) measurement system. First the light propagates through a dichroic beamsplitter and is focused ($f=10$ mm) into the $5$ $\mu$m-diameter-core fiber. Light that is transmitted through the 37 cm-long fiber can be detected on a power meter or imaged on an sCMOS camera as shown in Fig. 2(c). The power before the fiber can be measured with a power meter that flips into the beam path, which doubles as a beam block for background measurements. Any fluorescence generated in the core of the fiber and guided out in the direction opposite of the 810 nm pump beam is reflected at a dichroic beamsplitter and focused onto an Electron-Multiplying Charge-Coupled Device (EMCCD) camera as shown in Fig. 2(d). Spectral filters are used to remove scattered 810 nm light.
@@
Two fibers are used for the experiments: fiber 1 is used for the two lower sample concentration C2PA measurements (experiments 1 and 2) and fiber 2 is used for the highest sample concentration C2PA measurement (experiment 3) and for the E2PA measurement. The summed fiber scattering and toluene absorption coefficients are measured using long exposure images taken by a smartphone camera. The integrated intensity along the length of the fibers is fit to an exponential decay function as shown in Supporting Section 6. The coefficient at 810 nm is determined to be negligible whereas at 458 nm (near the fluorescence maximum) is 0.093 cm-1 for fiber 1 and 0.034 cm-1 for fiber 2. The measured transmission efficiency of the laser through the fiber was about $40\%$, $43\%$ and $47\%$ for experiments 1, 2 and 3 respectively; losses due to coupling, absorption in toluene ($11\%$) and scattering are all accounted for in this efficiency. This high transmission is consistent with the light primarily occupying the fundamental mode. For the SPDC, the transmission efficiency was considerably lower ($0.07\%$). This result is not entirely surprising since the SPDC beam occupies many spatial modes compared to the few-mode fiber. Furthermore, the fiber has a lower retention of light in higher-order modes and thus acts like a single-mode filter. Details on the measurement and definition of transmission efficiency for both laser and SPDC are given in Supporting Section 7.
The sample AF455 is chosen for our measurements due to its large C2PA cross section at 810 nm and its solubility in toluene. The “AF455” fluorophore is provided by the Air Force Research Laboratory43, 44. The toluene used to prepare the sample has a purity of $\geq 99.98\%$. The concentrations were calculated from the one-photon absorption spectra by use of a UV-VIS-NIR spectrophotometer (Agilent Cary 5000 Scan). The concentration of the sample is monitored over the course of a measurement and found to vary by $\leq 12\%$
4 Results and Discussion
@@
Classical two-photon excited fluorescence (C2PEF) datasets were gathered for three different concentrations of AF455 in toluene. Data series were acquired at a variety of laser powers until the uncertainty, set by the Allan deviation, reached a value at least 25% smaller than the photon rate derived from the measurement. For high excitation powers, the acquisition time was as little as five minutes, which produced uncertainty values of less than 1%. Our longest data series was acquired over a 13-hour period, which lowered the uncertainty to the point where the previously indistinguishable signal could be discerned.
@@
In Fig. 4(b) the fluorescence rate extracted from each frame’s background-subtracted image is plotted as a function of frame number for both raw (teal) and normalized (pink) data. The normalized data accounts for any changes in laser excitation power (see Supporting Section 8). In Fig. 4(c) the Allan deviation of the fluorescence rate is plotted for different numbers of frames ($N$) used for averaging for both raw (teal) and normalized (pink) data. For both Figs. 4(b) and (c), the overlap of the raw and normalized data is indicated by a purple trace. The Allan deviation should follow a $1/\sqrt{N}$ line in the absence of noise sources at that combined frame rate, thus this gold line is used to pick a near optimum number of frames to average. Here, 168 frames is selected, which corresponds to an Allan deviation of 1.3 cnt s-1. We average the data in portions of 168 frames and plot the result in Fig. 4(b) for the raw (gold) and normalized (blue) data. These measurements have an average value of 1.6 cnt s-1.
@@
The results of our fluorescence measurements for three concentrations of AF455 in toluene are plotted in Fig. 5(a), which shows the fluorescence count rate as a function of average excitation power. The fits to the datasets are linear regressions performed on a log-log scale. The slopes of the fits are all within $0.05$ of $2.00$, which confirms the two-photon origin of the signals. The lowest data point (acquired using the data shown in Fig. 4) is measured for the highest concentration at a power of $1.75$ nW. The signals shown here are compared with our model using measured and calculated parameters in Fig. 8 in Supporting Section 10 and show qualitative agreement.
@@
Using Eq. (9) of the Supporting Information, C2PA cross sections are derived using the fits to each of the three concentration datasets. We derive values of ($570\pm 190$) GM, ($340\pm 120$) GM and ($250\pm 80$) GM for the $2.30$ mM, $170\,\mu$M and $19.5\,\mu$M samples, respectively. The average of these is ($390\pm 80$) GM which is in line with the cross sections reported for AF45545, 16.
@@
The SPDC source was used exclusively with the $2.30$ mM sample. Using the high concentration sample maximizes the likelihood of measuring entangled two-photon excited fluorescence (E2PEF). The results from the SPDC data series are shown in Fig. 6. For this measurement the pump power was nearly maximum for an average of 49.2 mW, which generates $4.06\times 10^{11}$ photons s-1. Due to free space losses and fiber coupling losses, only $1.49\times 10^{8}$ photons s-1 make it into the fiber. The average background-subtracted image is shown in Fig. 6(a), where there is no discernible bright spot due to a signal. In Fig. 6(c), the Allan deviation is plotted, which is used to determine a near optimal averaging of 780 frames. This averaging corresponds to an uncertainty of 0.48 cnt s-1 for both the raw (teal) and normalized (pink) data. In Fig. 6(b), the fluorescence rate from each frame is plotted as well as the averaged (in 780 frame portions) fluorescence rate. The average fluorescence rate for the raw (gold) data is ($0.23\pm 0.48$) cnt s-1 and for the normalized (blue) data is ($0.22\pm 0.48$) cnt s-1. The magnitude of the signal is not significantly different from zero, and therefore we conclude that there is no resolvable signal.
Although E2PEF was indistinguishable from zero, there are numerous conclusions to be drawn from these measurements. The C2PEF dataset for the $2.30$ mM sample are plotted alongside our E2PEF measurement in Fig. 5(b). Here we show measured fluorescence count rates as a function of the average power of either the SPDC (orange) or laser (blue) source. For the E2PEF measurement, we couple about 36.5 pW of SPDC into the fiber. This plot demonstrates the difference between the SPDC power and the minimum laser power, which are only 48-fold apart.
We can compare these E2PA results with those done in free space (Table 1). Notably, the power of SPDC at the sample is 60-fold lower in the present experiment. This is due to the difficulties in coupling the SPDC into the few-mode fiber. Despite this lower power, the peak photon flux is $7.7\times 10^{19}$ photons cm-2s-1. In the free-space studies the SPDC peak photon flux was $2.1\times 10^{18}$ photons cm-2s-1, which is 37-fold lower. It should be noted that some of the advantage of the higher SPDC photon flux is hindered by the lower fraction of intact photon pairs propagating in the fiber, which is 0.00002% for fiber and 58% for free space. Here we have assumed that the losses are equivalent for both the signal and idler modes that are coupled into fiber. The low pair rate is accounted for in the calculations in Supporting Section 10.
Using Eq. (10) in Supporting Section 10, we set an upper bound on the E2PA cross section of this sample at $(4.0\pm 1.6)\times 10^{-21}$ cm2. This upper bound is for an entanglement time of 1070 fs and an entanglement area in the range $2.1$ $\mu$m2 to $18$ $\mu$m2. This entanglement time is an estimated value for the SPDC at the entrance of the fiber adapter. The broadening of the entanglement time in the fiber is among numerous parameters accounted for in the calculation of the upper bound (see Supporting Section 10).

### v2

@@ Abstract
We introduce a toluene-filled fiber platform for two-photon absorption measurements. By confining both the light and molecular sample inside the 5 $\mu$m hollow core of the fiber, we increase the distance over which the nonlinear light-matter interaction occurs. With only a 7.3 nL excitation volume, we measure classical two-photon absorption (C2PA) at an average laser power as low as 1.75 nW, which is a 45-fold improvement over a conventional free-space technique. We use this platform to attempt to measure entangled two-photon absorption (E2PA), a process with a limited regime where the quantum advantage is large. This regime arises due to a crossover from linear to quadratic scaling with photon flux as photon flux is increased. Recently, several teams of researchers have reported that E2PA cross-sections are much smaller than previously claimed. As a result, the linear scaling dominates at photon fluxes so low that it is extremely difficult or impossible to measure using conventional free-space techniques. In this report, we implement the first E2PA measurement using a waveguide. We see no evidence of E2PA, and we set an upper bound on the cross-section consistent with these recent reports.
keywords:
two-photon absorption, spontaneous parametric downconversion, liquid-core fiber, cross-section, low power, fluorescence detection
@@
Correlated photon pairs (red) are coupled into a fiber containing a solution of the molecule AF455. Fluorescence (blue) generated from the two-photon absorption process is guided out of the fiber.
Introduction
@@
Here, we show that with a toluene-filled fiber, 2PA cross-sections, which are the quantities that describe the strength of the 2PA process, can be measured at extremely low powers. To the best of our knowledge, this work is the first to show a 2PA measurement at an average power as low as 1.75 nW. This is a 45-fold lower power than that achieved in our previous results16 for low-power 2PA in free space for the same sample. These fiber and free-space measurements serve as a one-to-one comparison since both were designed to test the sensitivity limits of the platforms. We believe this is the first time that the advantage of the fiber platform for 2PA has been quantified.
@@
This sensitivity, paired with the ability to increase our laser power by up to nine orders of magnitude, demonstrates that the platform is well equipped to boost previously undetectable signals. This is necessary, for example, for measurements of C2PA cross-sections at much less efficient wavelengths or for much less efficient absorbers. Furthermore, rather than being limited to a high peak power laser17, unconventional light sources may produce sizeable 2PA signals.
@@
One such light source is an entangled photon pair source based on spontaneous parametric downconversion (SPDC). We employ an SPDC source in an attempt to measure entangled two-photon absorption (E2PA). To differentiate between E2PA and 2PA excited by a laser source, we refer to the latter process as classical two-photon absorption (C2PA). The E2PA process has been a subject of theoretical study since the 1980s (Refs. 18, 19, 20), driven by the idea that the strong temporal and spatial correlations of SPDC photons are ideal for 2PA. These temporal and spatial correlations are characterized by the entanglement time and area, respectively20. The entanglement time quantifies the width of the joint probability distribution of differences in arrival times of the two photons forming a pair, defined at the location of interest. The entanglement area is proportional to the width of the joint probability distribution of differences in positions of the two photons forming a pair, defined at the location of interest—to calculate entanglement area, the widths from the two transverse dimensions must be used.
The E2PA process is known to scale linearly with the excitation photon flux, rather than the quadratic scaling of C2PA, in the low-gain parametric downconversion (PDC) regime where, if it can be measured, it dominates over the quadratic component and thus demonstrates a quantum advantage. In the high-gain PDC regime, a quadratic scaling is recovered, with the possibility of up to three-fold signal boost relative to excitation with a coherent laser source. This boost reflects a higher likelihood of two photons overlapping due to the increased statistical fluctuations of the excitation source. A more extensive description of the SPDC regimes and expected quantum advantage for various bandwidth conditions of the pump, SPDC and absorption spectrum is covered in Refs. 21, 22.
@@
Recently, a wide range of evidence from a variety of experimental research studies has shown that molecular E2PA cross-sections are much smaller than previously reported. The majority of studies have reported no E2PA signal above the noise floor in the low-gain (Refs. 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 22) and moderate-gain PDC regimes (Refs. 16). One study reported an extremely weak linear-scaling signal (Ref. 33), however another study closely replicated this experiment and was unable to observe the reported signal22. Two studies have reported quadratic-scaling signals using PDC (Ref. 31, 22). In the report by Raymer and coworkers 22, a high-gain PDC source is used and the resulting signal exhibits a moderate boost relative to a coherent source, which is consistent with the increased photon number fluctuations of the excitation source. In the report by Wang and coworkers 31, a low-gain PDC source is used to excite upconversion nanoparticles that have long-lived intermediate states and thus can be excited via uncorrelated photon pairs despite using an excitation source with significant temporal separation between photon pairs. These experimental reports are reinforced by recent theoretical works that show that the linear scaling will dominate at photon fluxes so low that the resulting minuscule signal will be difficult or impossible to measure with current technology (Refs. 34, 35, 21, 36). Furthermore, the conditions under which the minuscule signal is maximized require careful selection of the SPDC source and two-photon absorber. These experimental and theoretical studies all consider conventional free-space excitation conditions.
Here we present, to the best of our knowledge, the first E2PA measurement implemented in a waveguide. Our approach uses a relatively high flux of photon pairs to maximize our likelihood of measuring a signal. In our setup, this high flux is generated by a moderate-gain PDC process. We do not observe E2PA and set an upper bound on the E2PA cross-section of the molecule AF455.
@@
In this paper, we discuss the principles of our experimental scheme and the design considerations along with our experimental implementation and its characterization. Next, we present our results for excitation of the molecule AF455 with both laser and SPDC photons and derive the C2PA and E2PA cross-sections, respectively. Finally we compare our results with those from previous free-space measurements16. More details on experimental characterization, data acquisition and data analysis are provided in the supplemental information.
Operating Principles
@@
The benefit of this scheme is illustrated in Fig. 1(a-b). In the free-space approach (Fig. 1(a)), light is focused into the cuvette to a waist $w_{0}$ and diverges on a length scale characterized by the Rayleigh range $z_{R}$ which scales quadratically with $w_{0}$. Thus, increasing beam intensity occurs at the cost of maintaining the intensity over a shorter distance $z_{R}$. For typical focusing conditions of $w_{0}\approx 30\,\mu$m for an 810 nm beam into toluene, $z_{R}\approx 5.2$ mm, which is about half of the cuvette’s length. In contrast, with the fiber-based approach (Fig. 1(b)), the beam waist $w_{0}$ is maintained with low loss over the entire propagation length of the fiber because of optical confinement. Typically, $w_{0}$ is much smaller in fiber than in free space. This high light intensity is advantageous for 2PA because two photons must be spatially overlapped at a molecule for absorption to occur. Maintaining the peak intensity of the focused beam over centimeter length-scales increases the excitation probability because of the interaction with a large number of molecules.
@@
The refractive indices as a function of wavelength for the toluene core and silica cladding are plotted in Fig. 1(c) using the known dispersion equations for the materials37, 38. The effective index of refraction of the fundamental and second-order modes are shown and are calculated using Ref. 39. We can estimate the number of modes that can propagate along the fiber, which has a core diameter $d$, using the V-number,
@@
Figure 2: Digital microscope images of one end of (a) the fiber and (b) the fiber inside of the tubing sleeve. (c) sCMOS image of the 810 nm excitation laser guided through the fiber. (d) EMCCD image of the fluorescence from AF455 guided out of the fiber. The last two images are taken in the image plane of the fiber face. Some features of the fiber and tubing sleeve are visible in both images as described in the Supporting Information.
In addition to toluene allowing broadband guidance of light in a silica fiber, this solvent also has a low absorption coefficient40 at the excitation and fluorescence wavelengths (0.0030 cm-1 and 0.0039 cm-1, respectively). This is a necessary condition otherwise the long length of the fiber will add little to no benefit since all the light would be absorbed after a short distance. For comparison, water has an absorption coefficient that is about one order of magnitude larger at 810 nm (0.0209 cm-140).
Methods
To prepare the fiber, we first cut clean facets on both ends to minimize light-coupling losses. We use a custom-built coil heater to remove about 20 mm of the polyamide coating from each end, then we cleave the fiber ends in the regions where the coatings were removed. We inspect the ends under a digital microscope (Keyence VHX 7000) to ensure the cuts are smooth and to check for particle contamination. Typical images are shown in Fig. 2(a)-(b). The fiber ends are placed in a tubing sleeve to prevent breakage and to allow for simple attachment to the custom-built fiber adapters (similar to those used in Refs. 41, 42, technical drawings provided in Ref. 43).
@@ Methods
The optical and microfluidic setup is illustrated in Fig. 3. Here we give a brief overview of the setup. A complete list of the components is given in Ref. 43. A tunable femtosecond pulsed laser operated at 810 nm (110 fs pulses, $\approx 9$ nm bandwidth) is used to generate SPDC photons for E2PA measurements and its direct output is used for C2PA measurements. For photon pair generation, the output is first directed through a second harmonic generation unit to produce 405 nm light ($\approx 3$ nm bandwidth). The average power is controlled using a half-wave plate and a polarizer. Spectral filters are used to remove any remaining 810 nm light or unwanted harmonics. A small portion of this beam is picked off by a beam sampler to monitor the power and beam pointing stability of the laser on a quadrant (quad) detector. The blue light is focused ($f=300$ mm lens) into a 10 mm-long type-0 periodically-poled potassium titanyl phosphate (ppKTP) crystal to generate photon pairs centered at 810 nm. The crystal is temperature controlled at $40.00\pm 0.01\,^{\circ}$. The photon pairs are collimated with a 50 mm lens and any remaining pump light is filtered out.
@@
In our previous report16 we found that this source has a spectral bandwith of approximately 75 nm full width at half maximum (FWHM) under these operating conditions. We measure that this source produces $8.25\times 10^{9}$ photons s-1 mW-1 using a scientific Complementary Metal–Oxide–Semiconductor (sCMOS) camera placed directly after the crystal and collimating lens. This corresponds to a mean photon number of $\approx 5000$ photons pulse-1 at our maximum pump power of 49.2 mW. We can estimate the number of SPDC spatial modes based on the fraction of the total photon rate that is collected into the few-mode fiber and accounting for losses due to fiber coupling, and absorption and scattering in toluene as discussed in the Supporting Information. This yields $\approx 740$ spatial modes, which serves as a lower bound since up to 16 spatial modes can propagate in our fiber. Using this estimate, we find that we generate $\approx 6.8$ photons pulse-1 spatial mode-1. The entanglement time of the SPDC is estimated using a discrete Fourier transform of the joint spectral amplitude, which is estimated using the measured joint spectral intensity (shown in Ref. 16) and a phase estimated using the calculated group delay dispersion accumulated by the SPDC between the center of the crystal and its position along the fiber length. The broadening of the entanglement time across the length of the fiber is taken into account in our calculations (Eq. (28) of Supporting Information). At the entrance of the fiber the entanglement time is 1070 fs. The entanglement area was not measured, but we estimate it to be in the range of 2.1 $\mu$m2 to 18 $\mu$m2. The lower bound is set by a diffraction limited spot size (see Ref. 16). The upper bound is set by the spot size of the fundamental mode of the fiber.
@@
For the C2PA measurements, the pulsed laser is directed around the nonlinear crystals and through a half-wave plate, polarizing beamsplitter and neutral density filters for control over the power. The light from both sources is recombined at a back-polished mirror. Either source is directed into the fiber-based two-photon excited fluorescence (2PEF) measurement system. First the light propagates through a dichroic beamsplitter and is focused ($f=10$ mm) into the $5$ $\mu$m-diameter-core fiber. Light that is transmitted through the 37 cm-long fiber can be detected on a power meter or imaged on an sCMOS camera as shown in Fig. 2(c). The power before the fiber can be measured with a power meter that flips into the beam path, which doubles as a beam block for background measurements. Any fluorescence generated in the core of the fiber and guided out in the direction opposite of the 810 nm excitation beam is reflected at a dichroic beamsplitter and focused onto an Electron-Multiplying Charge-Coupled Device (EMCCD) camera as shown in Fig. 2(d). Spectral filters are used to remove scattered 810 nm light.
@@
Two fibers are used for the experiments: fiber 1 is used for the two lower sample concentration C2PA measurements (experiments 1 and 2) and fiber 2 is used for the highest sample concentration C2PA measurement (experiment 3) and for the E2PA measurement. The summed fiber scattering and toluene absorption coefficients are measured using long exposure images taken by a smartphone camera. The integrated intensity along the length of the fibers is fit to an exponential decay function as shown in Fig. 7. The scattering coefficient at 810 nm is determined to be negligible whereas the absorption coefficient is equated to that found in literature40. At 458 nm (near the fluorescence maximum) the summed coefficient is measured to be 0.093 cm-1 for fiber 1 and 0.034 cm-1 for fiber 2.
The measured transmission efficiency of the laser through the fiber was about $40\%$, $43\%$ and $47\%$ for experiments 1, 2 and 3 respectively; losses due to coupling, absorption in toluene ($11\%$) and scattering are all accounted for in this efficiency. This high transmission is consistent with the light primarily occupying the fundamental mode. The average laser transmission efficiency serves as a best estimate for the transmission efficiency of the fundamental mode of SPDC ($43\%$). If we consider all spatial modes, including those that are irrelevant in the estimation of E2PA signals since they are not coupled into fiber, we can estimate an effective transmission efficiency of $0.07\%$. The huge difference in transmission efficiency between a single mode and $\approx 740$ modes illustrates that filtering the SPDC to a small number of modes may give the same result as the experiment we present. The fiber has a lower retention of light in higher-order modes and thus it effectively acts like a single-mode filter. Details on the definition and measurement of transmission efficiency for both laser and SPDC are given in the Supporting Information.
To estimate the efficiency of coupling one photon into fiber given that its spatially-correlated partner photon is coupled into fiber, we estimate an effective Klyshko44 efficiency using SPDCalc45. SPDCalc uses various parameters of our pump beam, crystal, lenses and fiber, to calculate the overlap integral of three Gaussian modes—one for signal photons, one for idler photons, and one for the collected single mode in fiber—along the length of the crystal. This yields $\eta_{K}^{{}^{\prime}}=0.94$. This differs from a measured Klyshko efficiency, $\eta_{K}$, because it does not account for any single photon loss between photon pair generation and collection into fiber. We multiply this value by the measured free-space transmission efficiency and the single-mode coupling efficiency to estimate $\eta_{K}=0.25$. This value is used to estimate that $25\%$ of photons coupled into fiber are part of an intact pair. See Eq. (Experimental Characterization) on Klyshko efficiency in the Supporting Information for more details.
@@
The sample AF455 is chosen for our measurements due to its large C2PA cross-section at 810 nm and its solubility in toluene. The “AF455” fluorophore is provided by the Air Force Research Laboratory46, 47. The toluene used to prepare the sample has a purity of $\geq 99.98\%$. The concentrations were calculated from the one-photon absorption spectra by use of a UV-VIS-NIR spectrophotometer (Agilent Cary 5000 Scan). The concentration of the sample is monitored over the course of a measurement and found to vary by $\leq 12\%$
Results and Discussion
Classical Two-Photon Absorption Measurements. Classical two-photon excited fluorescence (C2PEF) datasets were gathered for three different concentrations of AF455 in toluene. Data series were acquired at a variety of laser powers until the uncertainty, set by the Allan deviation, reached a value at least 25% smaller than the photon rate derived from the measurement. For high excitation powers, the acquisition time was as little as five minutes, which produced uncertainty values of less than 1%. Our longest data series was acquired over a 13-hour period, which lowered the uncertainty to the point where the previously indistinguishable signal could be discerned.
@@
In Fig. 4(b) the fluorescence rate extracted from each frame’s background-subtracted image is plotted as a function of frame number for both raw (teal) and normalized (pink) data. The normalized data accounts for any changes in laser excitation power (see Supporting Information). In Fig. 4(c) the Allan deviation of the fluorescence rate is plotted for different numbers of frames ($N$) used for averaging for both raw (teal) and normalized (pink) data. For both Figs. 4(b) and (c), the overlap of the raw and normalized data is indicated by a purple trace. The Allan deviation should follow a $1/\sqrt{N}$ line in the absence of noise sources at that combined frame rate, thus this gold line is used to pick a near optimum number of frames to average. Here, 168 frames is selected, which corresponds to an Allan deviation of 1.3 cnt s-1. We average the data in portions of 168 frames and plot the result in Fig. 4(b) for the raw (gold) and normalized (blue) data. These measurements have an average value of 1.6 cnt s-1.
The results of our fluorescence measurements for three concentrations of AF455 in toluene are plotted in Fig. 5(a), which shows the fluorescence count rate as a function of average excitation power. The fits to the datasets are linear regressions performed on a log-log scale. The slopes of the fits are all within $0.05$ of $2.00$, which confirms the two-photon origin of the signals. The lowest data point (acquired using the data shown in Fig. 4) is measured for the highest concentration at a power of $1.75$ nW. The signals shown here are compared with our model using measured and calculated parameters in Fig. 8 and show qualitative agreement.
@@
Using Eq. (Calculating a C2PA Cross-Section) of the Supporting Information, C2PA cross-sections are derived using the fits to each of the three concentration datasets. We derive values of ($570\pm 190$) GM, ($340\pm 120$) GM and ($250\pm 80$) GM for the $2.30$ mM, $170\,\mu$M and $19.5\,\mu$M samples, respectively. The average of these is ($390\pm 80$) GM which is in line with the cross-sections reported for AF45548, 16.
Entangled Two-Photon Absorption Measurement. The SPDC source was used exclusively with the $2.30$ mM sample. Using the high concentration sample maximizes the likelihood of measuring entangled two-photon excited fluorescence (E2PEF). The results from the SPDC data series are shown in Fig. 6. For this measurement the pump power was nearly maximum for an average of 49.2 mW, which generates $4.06\times 10^{11}$ photons s-1. A small number of modes of SPDC are effectively coupled into fiber, reducing the photon rate to $1.49\times 10^{8}$ photons s-1. The average background-subtracted image is shown in Fig. 6(a), where there is no discernible bright spot due to a signal. In Fig. 6(c), the Allan deviation is plotted, which is used to determine a near optimal averaging of 780 frames. This averaging corresponds to an uncertainty of 0.48 cnt s-1 for both the raw (teal) and normalized (pink) data. In Fig. 6(b), the fluorescence rate from each frame is plotted as well as the averaged (in 780 frame portions) fluorescence rate. The average fluorescence rate for the raw (gold) data is ($0.23\pm 0.48$) cnt s-1 and for the normalized (blue) data is ($0.22\pm 0.48$) cnt s-1. The magnitude of the signal is not significantly different from zero, and therefore we conclude that there is no resolvable signal.